OpenStudy (anonymous):

Factor the polynomial completely. x2 - 18

OpenStudy (kirbykirby):

difference of squares.. \(x^2-a^2=(x+a)(x-a)\)

OpenStudy (kirbykirby):

\((x+\sqrt{18})(x-\sqrt{18})=(x+3\sqrt{2})(x-3\sqrt{2})\)

OpenStudy (anonymous):

thanks!

OpenStudy (kirbykirby):

=]